If the product of two consecutive even integers is 34 larger than their sum then what are the two integers?: If the product of two consecutive even integers is 34 larger than their sum then what are the two integers?
Answer by ankor@dixie-net.com(4490)

f the product of two consecutive even integers is 34 larger than their sum then what are the two integers?
:
x(x+2) = x + (x+2) + 34
:
x^2 + 2x = 2x + 36
:
x^2 + 2x - 2x - 36 = 0
;
x^2 - 36 = 0
:
(x-6)(x+6)
;
x = +6, -6
:
Integers are 6 & 8, or -6, -4